OpenStudy (anonymous):

Don has been tracking the population of a town. The population was 7,220 last year and has been increasing by 1.45 times each year. How can Don develop a function to model this?

OpenStudy (anonymous):

A. \[ f(x)=(7220\times1.45)^x\] where 1.45 is the rate of growth

OpenStudy (anonymous):

B. f(x)=1.45(7220)^x where 7220 is the rate of growth

OpenStudy (anonymous):

C. f(x)=7220(1.45)^x where 1.45 is the rate of growth

OpenStudy (anonymous):

D. f(x)=7220(x)^(1.45) where 7220 is the rate of growth

OpenStudy (***[isuru]***):

hi, the answer should be.. \[f(x) = 7220\times(1.45)^{x}\] where 1.45 is the growth rate

OpenStudy (anonymous):

***[ISURU]*** is right, the answer is given in C)

OpenStudy (anonymous):

This question had me so confuzzled :/

OpenStudy (***[isuru]***):

ya need explanation or did ya got it now ?

OpenStudy (anonymous):

I got it :)